OpenStudy (anonymous):

the weight of man on the surface of earth is 588N. findd his maass taking g=98m/sec if the man were taken to moon,his weight would by 98N what is his mass on moon? caalculate acceleration due to gravity on the moon?

OpenStudy (anonymous):

can u solve..... plzz....!!!!!!

OpenStudy (anas.p):

calculate his mass first.... W=mg

OpenStudy (anas.p):

588=m*9.81

OpenStudy (anas.p):

m=588/9.8

OpenStudy (anas.p):

m=60

OpenStudy (anonymous):

wats the ans??

OpenStudy (anas.p):

mass will be same on moon and Earth. 60

OpenStudy (anas.p):

60kg

OpenStudy (anas.p):

On Moon W=mg 98/60=g g=1.63

OpenStudy (anonymous):

can u do it here?

OpenStudy (anonymous):

thoroly...?????

OpenStudy (anas.p):

First or second or both??

OpenStudy (anonymous):

both.....

OpenStudy (anas.p):

First we calculate his mass. The formula for weight is \[W=mg\] W is the weight(Force) m is mass g is the acceleration due to gravity. On earth g is always 9.81m/s. Your question tell us to take it as 9.8m/s So we take out the mass